A car tows a trailer and is restricted to pull a mass of not more than 500kg, the mass of each canned fish is 440g and the mass of each empty box is 27g. There are 15 boxes each contains 36 canned fish, determine the total mass in kg of each box



1
Expert's answer
2021-09-14T06:05:49-0400

Mass of the canned fish =440×36 =15,840g

Mass of the box =27g

To get the total Mass of each box ;

"\frac {(15,840+27)}{1000}kg"

Total mass in kg of each box=15.867Kg